If Joe has a $20.00 calling card. and long-distance calls cost $0.05 more per minute than local calls. If he made 35 minutes of local calls and 52 long distance calls, what was he paying per minute for each typy of call?
I am having problems with this...would someone explain it to me...thanks

Answer by checkley77(12844)

35x+52(x+.05)=20
35x+52x+2.60=20
87x=20-2.6o
87x=17.40
x=17.40/87
x=$.20 per minute for the local calls.
.20=.02=$.25 per minute for the long distance calls.
Proof:
35*.20+52*.25=20
7+13=20
20=20